# Modelling Lateral Spread in Wire Flat Rolling

### v1

@@ 4.2 Roll pressure
The roll pressure $\sigma^{(0)}_{yy}$ is found from (16) and plotted with respect to the contact angle, $\phi$ in Figure 5 for different reduction ratios. Results are compared with FE simulations from Vallellano et al. [16], the slab model from Kazeminezhad and Taheri [9] and the plane-strain slab model for the rolling of thin sheet in Erfanian et al. [17]. The same parameters as those used in Vallellano et al. [16] and described in the previous section are applied. The results for plane stress and plane strain are calculated from point A in Figure 2.
@@
Figure 5: Roll pressure from different methods; plane-strain slab method [17], current plane-stress method, and FE simulations [16] and Kazeminezhad and Taheri [9] slab method. Plotted for different reduction ratios against contact angle $\phi$. Other parameters used are $(\hat{\kappa},\hat{d}_{0},\hat{R},\mu)=(385/\sqrt{3}\,\mathrm{MPa},5\,\mathrm{mm},75\,\mathrm{mm},0.14)$.
@@
FE results for all reductions can be seen to have two distinct regions; a massive rise between the entrance and point A and a shallow rise from point A to the exit. This trend can be explained better when looking alongside Figure 6,
@@
Figure 6: A diagram of the deforming wire, seen from above (top figure) and from side (bottom figure). For symbols, refer to the main text. This is a modified version of Figure 3 from Carlsson [2].
@@
which schematically shows how the wire looks as seen from the side and above while rolling. The wire enters between rolls from the right-hand side with a circular cross-section and exits from the left with an almost rectangle cross-section, therefore, the contact surface forms rather like half of an ellipse. As explained by Carlsson [2] when the material starts to deform, surrounding parts of the wire are still in the elastic range and will therefore resist deformation. The situation may be compared to that of an indentation. As a result, a normal pressure will build up, resulting in a sharp rise in contact pressure soon after the entrance. As deformation continues, larger parts of the wire deform, reducing resistance from the remaining elastic regions, which in turn results in a pressure drop until point A. At this point, the material begins to flow laterally, and a typical friction hill develops, with a pressure peak forming between this point and the roll exit. [2].
@@
From the results, it can be seen that the current model correctly predicts the location of point A by assuming that the area of the cross-section at this point is the same as that of the wire before rolling. Both the current plane-stress model and Kazeminezhad and Taheri [9] correctly predict the roll pressure at point A, for larger reductions. From point A to the roll exit, the FE results lie between the plane-strain and plane-stress predictions, reflecting the presence of 3D material flow within the roll gap. Consequently, the results presented by Kazeminezhad and Taheri [9] show better agreement with the FE data. Another notable difference between the FE results and the perfect plastic models in Figure 5 is that the roll pressure starts at zero at the entrance and returns to zero at the exit. This behaviour arises from the inclusion of elasticity in the FE simulations, which is absent in the perfect plastic models.
@@ 5 Conclusion
This work develops a mathematical framework for wire rolling, emphasising the accurate prediction of lateral spread. The model with a rectangular cross-section is correlated to the real process by focusing on the region of the roll gap where lateral flow begins and assuming that, beyond this point, the cross-section in the real process can be approximated by a rectangular shape in the model with a matching area. Also, a general state of deformation is simplified to the plane-stress deformation with the assumption of a small width with respect to the length. The set of assumptions made leads to the simplest possible model, which is shown to successfully predict lateral spread across a range of wire diameters, reduction ratios, and different friction coefficients. Importantly, this model achieves accurate predictions without requiring fitting parameters or assuming additional factors. This holds true even though the assumption of plane-stress deformation is not strictly satisfied; comparison of roll pressure predictions with FE results suggests that the actual pressure distribution falls between plane-stress and plane-strain deformation. A possible explanation for this lies in the importance of the stress ratio, $s_{zz}/s_{yy}$, rather than the absolute stress values, in predicting the lateral spread. While the model may underestimate individual stresses, it appears to capture the correct ratio, leading to accurate predictions.
Nonetheless, having formulated stress and strain equations in all directions lays the foundation for a more general model. 3D numerical studies conducted using Abaqus software by Carlsson [2] and Vallellano et al. [16] show that the contact pressure distribution is more complex than the traditional pressure hill profile. Relaxing the assumption of a small rate of change in width in the current model would allow the problem to be modelled under a general stress state. Also, solving the higher order terms would enable accounting for the curvature of the edges.
Regarding the split observed during the rolling of brass wire, it is worth determining whether this issue stems from a manufacturing defect or if it represents a fundamental limitation of the rolling process for wires. Additionally, further studies are required to explore potential correlations between the challenges faced in experiments when achieving certain reductions in a single pass and the pressure singularity predicted by the current model.

### v2

@@ 4.3 Roll pressure
The roll pressure $\sigma^{(0)}_{yy}$ is found from (24) and plotted with respect to the contact angle, $\phi$ in Figure 6 for different reduction ratios.
@@
Figure 6: Roll pressure from different methods; plane-strain slab method (erfanian2025), current plane-stress method, and FE simulations (vallellano2008analysis) and kazeminezhad2006pressure slab method. Plotted for different reduction ratios against contact angle $\phi$. Other parameters used are $(\hat{\kappa},\hat{d}_{0},\hat{R},\mu)=(385/\sqrt{3}\,\mathrm{MPa},5\,\mathrm{mm},75\,\mathrm{mm},0.14)$.
Results are compared with FE simulations from vallellano2008analysis, the slab model from kazeminezhad2006pressure and the plane-strain slab model for the rolling of thin sheet in erfanian2025. The same parameters as those used in vallellano2008analysis and described in the previous section are applied. The results for plane stress and plane strain are calculated from point A in Figure 2.
@@
FE results for all reductions can be seen to have two distinct regions; a massive rise between the entrance and point A and a shallow rise from point A to the exit. This trend can be explained better when looking alongside Figure 7
@@
Figure 7: A diagram of the deforming wire, seen from above (top figure) and from side (bottom figure). For symbols, refer to the main text. Adapted from Figure 3 of carlsson1998contact.
@@
which schematically shows how the wire looks as seen from the side and above while rolling. The wire enters between rolls from the right-hand side with a circular cross-section and exits from the left with an almost rectangle cross-section, therefore, the contact surface forms rather like half of an ellipse. As explained by carlsson1998contact when the material starts to deform, surrounding parts of the wire are still in the elastic range and will therefore resist deformation. The situation may be compared to that of an indentation. As a result, a normal pressure will build up, resulting in a sharp rise in contact pressure soon after the entrance. As deformation continues, larger parts of the wire deform, reducing resistance from the remaining elastic regions, which in turn results in a pressure drop until point A. At this point, the material begins to flow laterally, and a typical friction hill develops, with a pressure peak forming between this point and the roll exit (carlsson1998contact).
@@
From the results, it can be seen that the current model correctly predicts the location of point A by assuming that the area of the cross-section at this point is the same as that of the wire before rolling. Both the current plane-stress model and kazeminezhad2006pressure correctly predict the roll pressure at point A, for larger reductions. From point A to the roll exit, the FE results lie between the plane-strain and plane-stress predictions, reflecting the presence of 3D material flow within the roll gap. Consequently, the results presented by kazeminezhad2006pressure show better agreement with the FE data. Yet, despite the model underestimates the pressure, it consistently provides accurate predictions of lateral spread. This may be attributed to the importance of the stress ratio $s_{zz}/s_{yy}$ — rather than the absolute magnitudes — in governing the lateral deformation of the wire. The model appears to capture this ratio correctly, despite discrepancies in individual stress components. Another notable difference between the FE results and the perfect plastic models in Figure 6 is that the roll pressure starts at zero at the entrance and returns to zero at the exit. This behaviour arises from the inclusion of elasticity in the FE simulations, which is absent in the perfect plastic models.
5 Parametric study
The present model, and indeed the governing equations behind it, depend only on three key non-dimensional parameters: the friction coefficient $\mu$, the aspect ratio $\delta$, and the reduction ratio $2\hat{h}_{1}/\hat{d}_{0}$ (recall $\hat{h}_{1}$ is the final half-thickness and $\hat{d}_{0}$ is the initial wire diameter). The aspect ratio $\delta=\hat{h}_{\!\>\!\mathrm{A}}/\hat{\ell}$ characterises the geometry of the deformation zone, and can also be expressed using (4) and (7) as
$\delta=\frac{\hat{h}_{\!\>\!\mathrm{A}}}{\hat{\ell}}=\frac{\sqrt{\pi}}{4}\!\left(\!\frac{\hat{R}}{\hat{d}_{0}}\!\left(\!\!\frac{\sqrt{\pi}}{2}-\frac{2\hat{h}_{1}}{\hat{d}_{0}}\!\right)-\frac{1}{4}\!\left(\!\!\frac{\sqrt{\pi}}{2}-\frac{2\hat{h}_{1}}{\hat{d}_{0}}\!\right)^{\!\!2}\,\right)^{\!\!-1/2}\!\!\!\!.$ (39)
Here, $\hat{R}/\hat{d}_{0}$ is also a non-dimensional parameter, and is arguably more intuitively connected to the dimensional rolling problem than is $\delta$. In what follows, therefore, we perform a parametric study by varying the three independent dimensionless parameters $\mu$, $2\hat{h}_{1}/\hat{d}_{0}$, and $\hat{R}/\hat{d}_{0}$, with $\delta$ then calculated using (39).
5.1 Range of validity
The range of validity of model at different reduction ratios and $\hat{R}/\hat{d}_{0}$ is shown for two friction coefficients in Figure 8.
Figure 8: The range of validity of model at different reduction ratios and $\hat{R}/\hat{d}_{0}$ for $\mu=0.1$ (left panel) and for $\mu=0.25$ (right panel). The shaded region represents the operating range of the model for which the normal stresses (equation (24)) are non-imaginary values. Below the red dotted line related to $\delta>0.3$ and the prediction should be interpreted with caution. The experimental data in the right panel are those reported in Figure 3.
For each combination of parameters, the valid range is determined by ensuring the computed normal stresses (equation (24)) remain real quantities (i.e., the expression under the square root remains positive). A more restrictive criterion for defining the range of validity of the model is to exclude large values of $\delta$, as dictated by the asymptotic formulation governing the solution accuracy. By way of indication, the line corresponding to $\delta=0.3$ is plotted in figure 8, with the region below this curve corresponding to $\delta>0.3$. Therefore, when the model is used to predict the width in cases where the roll and wire sizes are comparable and the reduction ratio is small, the results should be interpreted with caution, although some experimental data points presented in Figure 3 are located in $\delta>0.3$ region and the comparison in Figure 3 already shows good accuracy.
For a small friction coefficient, the model remains valid over a wider range of geometric parameters. In contrast, increasing the friction coefficient to $\mu=0.25$ considerably narrows the range of validity, as higher friction amplifies the $p^{(0)}$ term in equation (24).
Overall, the results show that the asymptotic model is most reliable for small aspect ratios (large $\hat{R}/\hat{d}_{0}$) and moderate reductions, which are representative of practical wire-rolling conditions. This is also consistent with the assumption of small $\delta$ in the asymptotic solution. It is worth noting that exceeding the max or min lines not only invalidates the present model (due to the square root going negative), but also correlates to a lack of obtainable result, as discussed in section 4.1; it may well be, therefore, that the physical rolling process undergoes a qualitative change beyond these parameters, and so these minima and maxima lines are dictated not by the approximations of the model but by the physics of the rolling process.
5.2 The effects of parameters on lateral spread
Figure 9: The effect of friction $\mu$ (left panel), reduction ratio $\hat{d}_{0}/2\hat{h}_{1}$ (middle panel) and roll size $\hat{R}/\hat{d}_{0}$ (right panel) on the predicted relative lateral spread $\hat{W}_{t}/\hat{d}_{0}$.
Figure 9 summarises a parametric study of the effect of different non-dimensional parameters on predicted lateral spread over the range of validity shown in Figure 8. The effect of the friction coefficient is shown in the left panel. Over the examined range $\mu\in[0.1,0.3]$, the width is only weakly sensitive to changes in friction for the low-reduction cases $\hat{d}_{0}/2\hat{h}_{1}=1.25$, for any roll sizes (blue and brown dashed lines). For the larger reduction case $\hat{d}_{0}/2\hat{h}_{1}=2$ the effect of friction is more pronounced (yellow and purple dashed lines): higher $\mu$ produces a noticeably larger lateral spread. Physically, this behaviour reflects that, at small reductions, material flow is only weakly controlled by tangential tractions, whereas at larger reductions friction increasingly impedes longitudinal exit flow and promotes lateral displacement, so $\mu$ becomes more influential.
As expected, all curves show a clear monotonic increase of lateral spread with increasing reduction ratio in the middle panel of Figure 9. However, the rate of increase is controlled by both $\mu$ and $\hat{R}/\hat{d}_{0}$. For the low-friction case ($\mu=0.10$, blue and brown curves), the width increases steadily but tends to flatten for large reductions, while for the higher-friction case ($\mu=0.30$, yellow and purple lines) the width rises much more rapidly with reduction. This demonstrates that friction amplifies the effect of reduction on lateral spread.
Finally, on the right panel of Figure 9, the influence of roll size $\hat{R}/\hat{d}_{0}$ depends on the reduction and friction. For the mild reduction case ($\hat{d}_{0}/2\hat{h}_{1}=1.25$, blue and brown lines), the relative width is almost insensitive to $\hat{R}/\hat{d}_{0}$ over the plotted range. By contrast, for the larger reduction ($\hat{d}_{0}/2\hat{h}_{1}=2$, yellow and purple lines), the width increases noticeably with $\hat{R}/\hat{d}_{0}$, and the increase is larger for higher friction ($\mu=0.3$, purple) than for low friction ($\mu=0.1$, yellow). A possible interpretation is that increasing $\hat{R}/\hat{d}_{0}$ increases the effective contact length and reduces local curvature effects, allowing more longitudinal material displacement to be converted into lateral spread for the higher-reduction cases; the conversion is enhanced when friction is larger.
@@ 6 Conclusion
This work develops a mathematical framework for wire rolling, emphasising the accurate prediction of lateral spread. A general state of deformation is simplified to the plane-stress deformation with the assumption of a small width with respect to the length. The set of assumptions made leads to the simplest possible first-principles model, which is shown to successfully predict lateral spread across a range of wire diameters, reduction ratios, and different friction coefficients. Importantly, this model achieves accurate predictions without requiring empirical fitting parameters or assuming additional factors. The validity of the model is demonstrated by comparison with both experimental and FE results, and the range of validity is defined. Interestingly, the range of validity also correlates with a lack of obtainable experimental result, suggesting an implication for the physics of the rolling process.
The assumption of plane-stress deformation is not strictly satisfied; comparison of roll pressure predictions with FE results suggests that the actual pressure distribution falls between that of plane-stress and plane-strain deformation. A possible explanation for this lies in the importance of the stress ratio, $s_{zz}/s_{yy}$, rather than the absolute stress values, in predicting the lateral spread; while the model may underestimate individual stresses, it appears to capture the correct ratio, leading to accurate predictions of lateral spread.
Nonetheless, having formulated stress and strain equations in all directions lays the foundation for a more general model. 3D numerical studies conducted using Abaqus software by carlsson1998contact and vallellano2008analysis show that the contact pressure distribution is more complex than the traditional pressure hill profile. Relaxing the assumption of a small rate of change in width in the current model would allow the problem to be modelled under a general stress state. Also, solving the higher order terms would enable accounting for the curvature of the edges.
Cold-rolled wire often exhibits a strong texture in the axial direction, while the texture in the cross-section can be quite different (carlsson1998contact). The von Mises function used to describe the yielding condition could be replaced with an anisotropic yield criterion (e.g., Hill’s 1948 yield criterion) with known anisotropy coefficients. In such a case, the solution structure would largely remain unchanged, except for the incorporation of weighted stresses. However, this scenario is of limited scientific significance because it is the rolling process itself which induces anisotropy. To accurately capture the anisotropy effect, it would be necessary for the anisotropy to evolve throughout the process. This would require a model describing how deformation induces anisotropy within the material, thereby adding an additional layer of complexity to the modelling process and is currently work in progress. Additionally, further studies are required to explore potential correlations between the challenges faced in experiments when achieving certain reductions in a single pass and the pressure singularity predicted by the current model.
@@ Acknowledgements
Appendix A Effect of initial cross-sectional shape
All results here have assumed that when a cylindrical wire is rolled, it first transforms into a square wire with aspect ratio $\hat{b}_{\!\>\!\mathrm{A}}/\hat{h}_{\!\>\!\mathrm{A}}=a=1$, as this preserves the symmetry of the cylindrical wire. In this appendix, we further investigate the effects of this assumption.
Figure 10: Effect of the assumed cross-sectional shape at location A on the predicted lateral spread. Different shapes corresponding to different values of $a=\hat{b}_{A}/\hat{h}_{A}$ are illustrated in the left panel.
Figure 10, computed as described in section 2.4 but for $b_{A}=a$, shows similar results to figure 4 but details the effect of varying $a$ on the results of our model. Figure 10 justifies that the choice of $a=1$ not only preserves symmetry but also gives the most accurate results compared with experimental results. Similar results were obtained for all experimental wire diameters, although Figure 10 plots the experimental results for the $\hat{d}_{0}=7.96\,\mathrm{mm}$ wire as these showed the greatest variation with $a$.